OpenStudy (lgbasallote):

by natural increase, a city whose population is 40,000 will double in 50 years. there is a net addition of 400 persons per yr because of the people leaving and moving unt the city. estimate the population in 10 yrs. hint: first find the natural growth proportionality factor. Ans: 50, 239 my answer is 49, 965. will post solution below

OpenStudy (lgbasallote):

double time formula: \[\frac{\ln 2}{k} = t\] \[\frac{\ln 2}{t} = k\] \[\frac{\ln 2}{50} = 0.0139\] \[\ln(\frac{x}{x_o}) = kt\] \[\frac{x}{x_o} = e^{kt}\] \[x = x_o e^kt\] \[x = 40,000 e^{0.0139 \times 10}\] \[x = 45, 965\] 400 people pr year so in 10 years there are 4,000 people increase so x = 49, 965

OpenStudy (lgbasallote):

so what did i do wrong

ganeshie8 (ganeshie8):

400 * 10 = 4000 this cannot be linear growth it grows naturally after 10 years it sums to : 400e^9k + 400e^8k +..... 400e^k + 400

OpenStudy (lgbasallote):

so im going to solve that one by one?

ganeshie8 (ganeshie8):

geometric sequence r = e^-k n = 10

OpenStudy (lgbasallote):

oh lol of course

OpenStudy (lgbasallote):

remind me what the formula is?

ganeshie8 (ganeshie8):

i just googled : http://regentsprep.org/Regents/math/algtrig/ATP2/GeoSeq.htm

OpenStudy (lgbasallote):

nice thanks

OpenStudy (lgbasallote):

why is r = e^-k?

ganeshie8 (ganeshie8):

400e^9k + 400e^8k +..... 400e^k + 400 common ratio : 2nd term/1st term

OpenStudy (lgbasallote):

oh of course...

ganeshie8 (ganeshie8):

im getting that answer... maybe k has to be more significant.. i mean upto 5 or 6 decimal placces

OpenStudy (lgbasallote):

\[s_{10} = \frac{a_1 (1-r^n)}{1-r} \implies \frac{400(1-e^{-0.0139 \times 10})}{1- e^{-0.0139}}\] is that right?

ganeshie8 (ganeshie8):

right

ganeshie8 (ganeshie8):

hmmm no wait

OpenStudy (lgbasallote):

no?

ganeshie8 (ganeshie8):

a = 400e^9k ...

OpenStudy (lgbasallote):

hmm okay \[\large S_{10} = \frac{400e^{9 \times 0.0139} (1 - e^{-0.0139 \times 10})}{1 - e^{-0.0139}}\] correct?

ganeshie8 (ganeshie8):

yh looks correct

OpenStudy (lgbasallote):

i got 50, 226

ganeshie8 (ganeshie8):

ok.. put the exact value of k in caculator.. maybe we will get the answer given

ganeshie8 (ganeshie8):

k = ln2/50

ganeshie8 (ganeshie8):

good luck :)

OpenStudy (lgbasallote):

i tried more exact and i got 50, 210

ganeshie8 (ganeshie8):

hmmm i think i give up... not getting the book answer :\

OpenStudy (lgbasallote):

lol...i think it's a matter of significant figures...how many do you think i should use

ganeshie8 (ganeshie8):

nope.. i have put k = ln2/50 still not getting the book answer

OpenStudy (anonymous):

\[40,000\times 2^{\frac{1}{5}}+400\times 10=49,948\] rounded

OpenStudy (lgbasallote):

@satellite73 yes that is what i got too at first but the answer should be 50,239. any idea how that happened?

OpenStudy (anonymous):

no

OpenStudy (lgbasallote):

aww okay :(